# Dance-to-Music Generation with Encoder-based Textual Inversion of Diffusion Models

### v1

@@ 3.2 Rhythm Encoder
We combine traditional feature extraction with a Multilayer Perceptron (MLP). “Dance to the rhythm” serves as an important connection point for synchronizing music and video. Conversely, “drop the beat to the motion” is crucial for the task of creating music for dance videos. During dancing, dancers typically perform actions or transitions in sync with specific musical beats. The initiation and transitions of movements align with points of local maximum acceleration in kinematics. Building upon LORIS [51], we calculate the first-order difference of the 2D keypoints $p(t,j,c)$ of the dancers over time to obtain the motion velocities. We use $v_{x}$ and $v_{y}$ to represent the velocities in the x and y directions, respectively. Subsequently, we discretize them into $K$ intervals, a process which we refer to as direction discretization:
$v_{Q}(t,j,k)=Vl_{\theta}(t,j,k),$ (4)
$l_{\theta}(t,j,k)=\begin{cases}1,&\text{if }\text{$k=[\frac{\theta}{{2\pi}/{K}}]$}\\ 0,&\text{otherwise}\\ \end{cases}$ (5)
$V=\sqrt{{v_{x}}^{2}+{v_{y}}^{2}},\\ \theta=\arctan\left(\frac{v_{y}}{v_{x}}\right),$ (6)
@@
where $\theta$ represents the direction angle, and $V$ represents the magnitude of the velocity. Continuing with the calculation of the first-order difference of $v_{Q}$, we obtain the discrete acceleration $a_{Q}$.
@@
We retain the positive values of acceleration and then calculate the sum of the accelerations in the joint dimensions and directional dimensions to obtain the total acceleration, denoted as $a$:
$a=\sum_{j,k}a_{Q}(t,j,k).$ (7)
@@
Now we determine the local maxima of $a$ within a time window. Then, we create a rhythm sequence $r$ where positions corresponding to the local maxima are set to 1, and all others are set to 0. We created a projector to map $r$ to the textual space, resulting in an embedding denoted as $v_{i*}$, which replaces the text embedding for the pseudo-word “*”.
3.3 Genre Encoder
@@
Reconstructing the audio mel-spectrogram solely from rhythmic information may introduce music genre information into the rhythmic pseudo-word, which hinders attribute disentanglement. Furthermore, it also results in a lack of control in music generation based on dance genre. To overcome this, we introduce a genre encoder to enhance the diffusion model’s reconstruction capability by incorporating additional information. We represent dance genres using one-hot encoding. By employing a combination of linear and activation layers, we map the one-hot encoding to the textual space, effectively replacing the text embedding associated with the pseudo-word “@” with the corresponding genre embedding $v_{i@}$.
4 Experiments
@@
Table 1: Qualitative comparison with state-of-the-arts methods. The best results are in bold and the second best results are underlined.
Rhythm Audio Quality Genre MOS
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ FAD $\downarrow$ Genre KLD $\downarrow$ Coherence $\uparrow$ Quality $\uparrow$
Ground Truth - - - - - 4.29 4.23
CMT [9] 0.3368 0.1515 0.2090 16.54 0.7297 1.81 3.25
CDCD [55] 0.4233 0.2151 0.2852 16.47 0.4592 1.96 1.41
LORIS [51] 0.2603 0.1073 0.1520 16.81 0.4405 1.66 1.33
MDM [45] 0.3798 0.4185 0.3982 4.812 0.5878 2.95 2.78
Ours (MUSICGEN) 0.4118 0.3874 0.3992 6.014 0.3354 3.52 3.59
Ours (Riffusion) 0.4761 0.4398 0.4572 3.416 0.3878 3.10 3.17
Figure 3: Qualitative examples of beat correspondence. The time scale interval is 0.1s. Our analysis of beats in real music and generated results from two frameworks show that the majority of the generated beats closely match the ground truth, with offsets below 0.2s. The Riffusion-based framework demonstrate slightly better beat alignment compared to MUSICGEN-based framework, consistent with quantitative metrics.
4.1 Experimental Setup
@@
In our experimental study, we utilize the primary perspective data from the AIST++ dance motion dataset [25]. This dataset comprises 70 songs paired with 460 Choreographies, evenly distributed across 10 different dance genres, along with corresponding 2D keypoints. We employ the official training, validation, and testing sets and segment the data into 5.12-second clips. As a result, our final training, validation, and testing sets consist of 2744, 36, and 36 samples, respectively. Furthermore, we transform the audio into corresponding mel-spectrograms to be used for experiments conducted on the Riffusion model. We utilize mmPose [7] to obtain 2D skeletons (dance motion sequences) of in-the-wild data from YouTube, using HRNet [44] pre-trained on the MS COCO [27] dataset.
@@
We conduct experiments on our encoder-based textual inversion method using two prominent music generation frameworks: diffusion and transformer. The base models employed for these frameworks were Riffusion [11] and MUSICGEN [8], respectively. For Riffusion, we fix the parameters of the Latent Diffusion Models (LDMs) [36] and the text encoder, except for the encoder branches. We use the default hyperparameters of LDMs and employ a base learning rate of 0.0005. The model training is executed on an NVIDIA GeForce RTX 3090 GPU with a batch size of 1, taking approximately 12 hours to complete over 50 epochs. As for MUSICGEN, we fix the parameters of the audio compression model, transformer-based audio token decoder, and T5 encoder. We apply the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.95$, and a weight decay of 0.01. A warm-up learning rate of 1e-4 is implemented for all layers during the initial 6000 training iterations. Both encoders are jointly optimized. Training the model for 20 epochs using a single NVIDIA A40 GPU with a batch size of 1 requires approximately 10 hours.
4.2 Evaluation Metrics
@@
Regarding the musical beats, a tolerance offset of 1 second is easily discernible to the human ear. To enhance the accuracy of evaluation, we refine the rhythm evaluation metrics by adjusting the tolerance offset from 1 second to 0.2 second. Specifically, we follow the setup of LORIS [51], where $B_{g}$ denotes the number of musical beats in the generated music, $B_{t}$ denotes the number of musical beats in the ground-truth music, and $B_{a}$ denotes the number of aligned musical beats. The evaluation metrics related to rhythm alignment are as follows:
  • Beat Coverage Score (BCS): the ratio of aligned musical beats to the generated musical beats ($B_{a}/B_{g}$);
  • Beat Hit Score (BHS): the ratio of aligned musical beats to the ground truth beats ($B_{a}/B_{t}$);
@@
Frechet Audio Distance (FAD) [24] is widely used for evaluating audio quality by measuring the similarity between the generated audio and the ground truth. A lower FAD value indicates more close to the ground truth. We report the FAD based on VGGish audio embedding model [17], which is pre-trained on the YouTube-8M audio dataset [2].
Genre
We employ MS-SincResNet [6], a pre-trained genre classifier, to compute the genre Kullback Leibler Divergence (KLD) score. The KLD score measures the distance between the predicted genre probabilities of the generated music and the ground-truth genre labels, thereby evaluating the genre attributes of the generated music.
Table 2: Ablation study of our method. “CLAP” represents utilization of CLAP audio embeddings to replace genre labels. “MLP” represents the adoption of a Multi-Layer Perceptron (MLP) as the projector in the rhythm encoder. “Attn+Pos” denotes the employment of attention mechanism and positional encoding as the projector in the rhythm encoder.
Rhythm Audio Quality Genre
[…]
Ours (MUSICGEN) CLAP 0.3925 0.3155 0.3498 6.821 0.3593
MLP+Genre label 0.3958 0.3643 0.3794 6.469 0.3283
Attn+Pos 0.4118 0.3874 0.3992 6.014 0.3353
Ours (Riffusion) CLAP 0.4480 0.3085 0.3654 4.611 0.7053
MLP+Genre lable 0.4761 0.4398 0.4572 3.416 0.3878
Attn+Pos 0.3806 0.2768 0.3205 3.814  0.3821
4.3 Quantitative Evaluation
We conducted experiments by deploying our method in two mainstream frameworks: the diffusion-based Riffusion [11] and the transformer-based MUSICGEN [8]. Additionally, we compared our approach against four state-of-the-arts methods: CMT [9], CDCD  [55], LORIS [51], and MDM  [45]. Specifically, CMT is based on the transformer architecture, while the remaining methods utilize diffusion model. We re-implemented CDCD and LORIS using our dataset and evaluated CMT and MDM using the models provided by the original authors. Our experimental results based on two frameworks and their comparisons with other methods are shown in Table  1. Our experiments based on Riffusion achieve optimal results in terms of rhythm and audio quality while ranking slightly lower in genre performance. Experiments based on MUSICGEN demonstrate rhythm performance comparable to that of MDM and achieve optimal results in genre performance, but slightly weaker audio quality. The MIDI-based approach, CMT, performs significantly worse than our method across all metrics. Both CDCD and LORIS exhibit poor performance in terms of rhythm and audio quality. Specifically, CDCD demonstrates a decent BCS score but performs poorly in terms of BHS, indicating its inclination towards generating fewer yet accurate beats.
User study
@@ 4.3 Quantitative Evaluation
To subjectively evaluate the performance of our method, we conduct a Mean Opinion Score (MOS) test. We provide participants with guidelines outlining the evaluation criteria for dance-to-music generation prior to the test. During the test, 78 participants rate a total of 70 generated samples, including 10 samples from each method and their corresponding ground truth. The evaluations are based on two criteria: dance-music coherence and audio quality. Dance-music coherence refers to the degree of alignment between the music and the dance video in terms of rhythm and style. Audio quality refers to the musicality of the audio (e.g., noise level, richness of melody). The rating scale ranges from 1 (best) to 5 (worst). As shown in Table 1, our approach based on MUSICGEN outperforms other methods in both metrics and achieves an acceptable margin compared to the ground truth. Interestingly, the results from MUSICGEN exhibit the best performance in dance-music coherence, indicating a tendency to generate music that closely align with human perception.
4.4 Qualitative Results
We visualized the beats of the generated music and compared them with the beats of ground truth, as shown in Fig. 3. Our method consistently generates reliable beats with offsets below 0.2, and the Riffusion framework shows superior beat alignment compared to MUSICGEN.
@@ 4.4 Qualitative Results
The generated audio samples, showcasing the comparison between our method and other approaches, can be accessed on the static webpage provided within the supplementary materials. The samples indicate that our model can generate more plausible dance music by guiding text-to-music models with prompts containing pseudo-words. Furthermore, our model demonstrates the capability to generate plausible beat-aligned dance music on in-the-wild data (e.g., film clips, virtual human dance videos, amateur dance videos). Furthermore, our method performs well in various other physical activities. Benefiting from the editability of the text-to-music model, our method exhibits the ability to generate diverse music that harmonizes with the dance video. The generated music aligns with the emotional atmosphere depicted in the videos, showcasing the robustness of our approach and its potential for real-world applications.
4.5 Ablation Study

### v2

@@ 3.2. Rhythm Encoder
We combine traditional feature extraction with a projector. “Dance to the rhythm” serves as an important connection point for synchronizing music and video. Conversely, “drop the beat to the motion” is crucial for the task of creating music for dance videos. Dancers typically perform actions or transitions in sync with specific musical beats during dancing. The initiation and transitions of movements align with points of local maximum acceleration in kinematics. Based on LORIS (Yu et al., 2023), we calculate the first-order difference of the 2D keypoints $p(t,j,c)$ of the dancers over time to obtain the motion velocities, where $j$ represents joints and $c$ denotes coordinates. We use $v_{x}$ and $v_{y}$ to represent the velocities in the x and y directions, respectively. Subsequently, we discretize them into $K$ intervals, a process which we refer to as direction discretization:
(4) $v_{Q}(t,j,k)=Vl_{\theta}(t,j,k),$
(5) $l_{\theta}(t,j,k)=\begin{cases}1,&\text{if }\text{$k=[\frac{\theta}{{2\pi}/{K}}]$}\\ 0,&\text{otherwise}\\ \end{cases}$
(6) $V=\sqrt{{v_{x}}^{2}+{v_{y}}^{2}},\\ \theta=\arctan\left(\frac{v_{y}}{v_{x}}\right),$
@@
where $\theta$ represents the direction angle, and $V$ represents the magnitude of the velocity. By calculating the first-order difference in the temporal dimension of $v_{Q}$ and aggregating the acceleration values across bins, we obtain the discrete acceleration $a_{Q}$.
@@
We retain the positive acceleration values and compute the sum of these accelerations across joint and directional dimensions to obtain the total acceleration, denoted as $a$:
(7) $a=\sum_{j,k}a_{Q}(t,j,k).$
@@
Next, we determine the local maxima of $a$ within a given time window. Subsequently, we construct a rhythm sequence $r$, where positions corresponding to the local maxima are set to 1, while all other positions are set to 0. We then design a projector to map $r$ to the textual space, resulting in an embedding denoted as $v_{i*}$, which replaces the text embedding for the pseudo-word “*”.
3.3. Genre Encoder
@@
Reconstructing the audio mel-spectrogram solely from rhythmic information may introduce music genre information into the rhythmic pseudo-word, which hinders attribute disentanglement. Furthermore, it also results in a lack of control in music generation based on the dance genre. To overcome this, we introduce a genre encoder to enhance the generative model’s reconstruction capability by incorporating genre information. We represent dance genres using one-hot encoding. By employing a combination of linear and activation layers, we map the one-hot encoding to the textual space, effectively replacing the text embedding associated with the pseudo-word “@” with the corresponding genre embedding $v_{i@}$.
4. Experiments
@@
Table 1. Quantitative evaluation and user study results in comparison with state-of-the-arts methods on AIST++ dataset. The best results are in highlighted bold and the second best ones are underlined (same in the following tables).
Rhythm Quality Genre MOS Inference Time
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ TD $\downarrow$ FAD $\downarrow$ CLAP $\uparrow$ Coherence $\uparrow$ Quality $\uparrow$ s/clip $\downarrow$
Ground Truth - - - - - - 4.26 4.25 -
CMT (Di et al., 2021) 0.3368 0.1515 0.2090 21.74 16.54 0.4454 1.79 3.08 6.44
CDCD (Zhu et al., 2022b) 0.4233 0.2151 0.2852 19.25 16.47 0.3032 2.02 1.35 5.72
LORIS (Yu et al., 2023) 0.3721 0.3371 0.3537 17.80 13.15 0.6180 2.56 2.42 21.2
MDM (Tan et al., 2023) 0.3798 0.4185 0.3982 22.96 4.812 0.5793 2.97 2.58 3.71
Ours (AudioLDM) 0.4419 0.3605 0.3971 22.73 8.522 0.7030 2.95 3.02 14.82
Ours (MUSICGEN) 0.4118 0.3874 0.3992 16.06 6.014 0.4685 3.56 3.54 11.48
Ours (Riffusion) 0.4761 0.4398 0.4572 20.34 3.416 0.7680 3.24 3.15 8.49
Table 2. Comparision results of the phase-aligned version of each output. OA, OM, and OR represent Ours (AudioLDM), Ours (MUSICGEN), and Ours (Riffusion), respectively.
Rhythm Quality Genre
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ TD $\downarrow$ FAD $\downarrow$ CLAP $\uparrow$
CMT 0.4038 0.1850 0.2538 21.74 16.94 0.5135
CDCD 0.4619 0.2318 0.3087 21.66 16.71 0.2845
LORIS 0.4476 0.3438 0.3889 21.39 13.42 0.6216
MDM 0.3460 0.4270 0.3823 21.44 4.258 0.5889
OA 0.4073 0.3351 0.3677 26.40 8.739 0.6853
OM 0.4368 0.4341 0.4354 15.19 5.762 0.6384
OR 0.4452 0.4251 0.4349 20.34 3.201 0.7453
4.1. Experimental Setup
@@
We evaluate our method on two dance-music datasets, i.e., AIST++ (Li et al., 2021) and our newly collected InDV dataset, both consisting of ten genres. The AIST++ dataset comprises 70 songs paired with 460 Choreographies, evenly distributed across 10 different dance genres, along with corresponding 2D keypoints. We employ the official training, validation, and testing sets and segment the data into 5.12-second clips. As a result, our final training, validation and testing sets consist of 2744, 36, and 36 samples, respectively. It is free for research purpose and we use this dataset for the main experiments and evaluations. We also collect and annotate a new dance-music dataset called InDV (In-the-wild Dance Videos), which contains 595 5.12-second clips with 216 songs. The dataset is categorized into 10 genres: Ballet, Breaking, Chinese Traditional Dance, Hip-hop, Jazz, Latin, Locking, Popping, and Waacking. The training, validation, and testing sets consist of 544, 25, and 26 samples, respectively. We employ HRNet (Sun et al., 2019) in mmPose (Contributors, 2020) to obtain 2D skeletons (dance motion sequences) from in-the-wild data. The video fps for the keypoints is 60.
@@
We conduct experiments on our encoder-based textual inversion method using three prominent music generation frameworks: diffusion-based models Riffusion (Forsgren and Martiros, 2022) and AudioLDM (Liu et al., 2023), and autoregressive model MUSICGEN (Copet et al., 2023). For Riffusion, we use the default hyperparameters of LDMs and employ a base learning rate of 0.0005. The model training is executed on an NVIDIA GeForce RTX 3090 GPU, taking approximately 12 hours to complete over 50 epochs. As for MUSICGEN, we apply the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.95$, and a weight decay of $0.01$. A warm-up learning rate of 1e-4 is used for all layers during the initial $6000$ training iterations. We train the network for 20 epochs on a single NVIDIA A40 GPU, taking approximately 10 hours. For AudioLDM, we keep the default hyperparameters and train it for 50 epoch on a single NVIDIA A40 GPU, taking approximately 11 hours.
4.2. Evaluation Metrics
@@
Regarding the musical beats, a tolerance offset of one second is easily discernible to the human ear. To enhance the accuracy of evaluation, we adjust the tolerance offset from 1 second to 0.2 second. Specifically, we follow the setup of LORIS (Yu et al., 2023), where $B_{g}$ denotes the number of musical beats in the generated music, $B_{t}$ denotes the number of musical beats in the ground-truth music, and $B_{a}$ denotes the number of aligned musical beats. We also conduct evaluation on global rhythm metrics. The evaluation metrics related to rhythm alignment are as follows:
  • Beat Coverage Score (BCS): the ratio of aligned musical beats to the generated musical beats ($B_{a}/B_{g}$).
  • Beat Hit Score (BHS): the ratio of aligned musical beats to the ground truth beats ($B_{a}/B_{t}$).
@@
  • Tempo Difference (TD): average L1 norm of tempo difference between generated and ground truth music.
@@
Frechet Audio Distance (FAD) (Kilgour et al., 2019) is widely used for evaluating audio quality by measuring the similarity between the generated audio and the ground truth. A lower FAD value indicates more close to the ground truth. We report the FAD based on VGGish audio embedding model (Hershey et al., 2017), which is pre-trained on the YouTube-8M audio dataset (Abu-El-Haija et al., 2016).
Genre similarity
We employ CLAP (Wu et al., 2023a), a pre-trained large-scale contrastive language-audio model, to compute genre similarity. The CLAP score evaluates the degree of similarity between the CLAP embeddings of the generated audios and the ground truth audios.
Figure 3. Music visualization examples. “GT” and “Ours” represents “Ground Truth” and “Ours (MUSICGEN)”, respectively.
4.3. Quantitative Evaluation
AIST++ dataset (Li et al., 2021)
We conducted experiments on AIST++ dataset by deploying our method on three mainstream base models: the diffusion-based Riffusion (Forsgren and Martiros, 2022), AudioLDM (Liu et al., 2023) and the autoregressive MUSICGEN (Copet et al., 2023). Additionally, we compared our approach against four state-of-the-arts methods: CMT (Di et al., 2021), CDCD (Zhu et al., 2022b), LORIS (Yu et al., 2023), and MDM (Tan et al., 2023). We re-implemented CDCD and LORIS and evaluated CMT and MDM using the models provided by the original authors. Our experimental results on three base models and their comparisons with other methods are shown in Table 1. Our method outperforms others across all metrics. Specifically, Riffusion-based experiments achieve optimal results on all objective metrics except TD, while MUSICGEN-based experiments achieve the best results in subjective metrics. The utilization of large-scale pre-trained models leads to longer inference times. We additionally compare the phase-aligned versions of each output (see Table 2), and our method achieves the optimal overall rhythm alignment (F1-score), indicating that our method achieves local adaptation. The slight weakness of BCS might be due to CMT, CDCD and LORIS tending to generate fewer beats. While global alignment easily improves their beat precision (BCS), the recall rate (BHS) still shows a significant weakness. In terms of genre similarity and audio quality, our method maintains an obvious advantage. The rhythm alignment metrics for certain superior methods decline because the global offset disrupts the local alignment of the generated results to some extent.
Figure 4. Qualitative examples of beat alignment. The time scale interval is 0.1s. The distribution of beats show that the majority of the generated beats closely match the ground truth, with offsets below 0.2s. The Riffusion-based model demonstrates slightly better beat alignment compared to MUSICGEN-based model, consistent with quantitative metrics.
Table 3. Quantitative evaluation on InDV dataset. OM and OR represent Ours (MUSICGEN) and Ours (Riffusion), respectively.
Rhythm Quality Genre
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ TD $\downarrow$ FAD $\downarrow$ CLAP $\uparrow$
CMT 0.3513 0.2045 0.2585 19.80 12.65 0.4519
CDCD 0.3155 0.1905 0.2375 22.80 15.77 0.0747
LORIS 0.2782 0.2624 0.2701 26.15 13.51 0.2975
MDM 0.2439 0.1985 0.2189 21.98 11.30 0.4404
OM 0.3594 0.4012 0.3792 18.80 5.791 0.4048
OR 0.3613 0.4562 0.4032 12.57 5.633 0.5161
InDV dataset
Compared to AIST++, our InDV dataset is a more challenging dataset with in-the-wild video settings that contains diverse dance movements. Table 3 shows the results of the quantitative evaluation of the experiments in the InDV dataset. Our method outperforms other approaches across multiple metrics, which demonstrates the overall robustness of our approach.
Table 4. Comparison results with music generated at the target tempo.
Rhythm Quality Genre
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ FAD $\downarrow$ CLAP $\uparrow$
Mustango 0.3849 0.3247 0.3522 13.33 0.4566
Phase-aligned Mustango 0.3964 0.3260 0.3578 14.08 0.5230
Ours (MUSICGEN) 0.4118 0.3874 0.3992 6.014 0.4685
Ours (Riffusion) 0.4761 0.4398 0.4572 3.416 0.7680
Comparison with music generated at the target tempo
We use Mustango (Melechovsky et al., 2024) to generate music at the same tempo as the groundtruth and perform phase alignment. The comparison results are shown in Table 4. Our method significantly outperforms the results of Mustango, indicating that our approach provides local adaptation compared to the global control of the tempo.
User study I
@@ 4.3. Quantitative Evaluation
We conduct a Mean Opinion Score (MOS) test. We provide participants with guidelines outlining the evaluation criteria for dance-to-music generation prior to the test. During the test, 82 participants rate a total of 80 samples, including 10 samples from each method and their corresponding ground truth. The evaluations are based on two criteria: dance-music coherence and audio quality. Dance-music coherence refers to the degree of alignment between the music and the dance video in terms of rhythm and style. Audio quality refers to the musicality of the audio (e.g., noise level, richness of melody). The rating scale ranges from 1 (best) to 5 (worst). As shown in Table 1, our approach based on MUSICGEN outperforms other methods in both metrics and achieves an acceptable margin compared to the ground truth.
Figure 5. Visualization of results for videos with changes in tempo. The frequency of audio peaks changes in accordance with the variations in video tempo and aligns with dance movements. The first row illustrates an example of slowed-down tempo, while the second row showcases an example of accelerated tempo.
4.4. Qualitative Results
We visualize the waveforms and corresponding dance movements for both the ground truth and the generated music (see Fig. 3). Our results show richer dynamics that closely resemble the ground truth, while maintaining alignment with the dance movements. We also compare the beats of the generated music with those of the ground truth (see Fig. 4), demonstrating consistent generation of reliable beats with offsets below 0.2. The generated audio samples, showcasing the comparison between our method and other approaches, can be accessed on the static webpage provided within the supplementary materials. The samples show that our method generates music with superior coherence and higher quality compared to other approaches.
To verify that our method can adapt to changes in tempo, we manually change the speed of dance videos. The generated audio can still be aligned with the video rhythm while exhibiting tempo variations, as shown in Fig. 5. Detailed samples can be found in the supplementary video.
@@ 4.4. Qualitative Results
Furthermore, samples in our supplementary materials also demonstrate the ability of our method to generate plausible beat-aligned dance music on in-the-wild data (e.g., film clips, virtual human dance videos, and amateur dance videos). Moreover, our method performs well in various other physical activities. The generated music aligns with the emotional atmosphere depicted in the videos, showcasing the robustness of our approach and its potential for real-world applications. Benefiting from the editability of the text-to-music model, our method demonstrates the ability to generate diverse music that aligns with the same dance videos.
4.5. Ablation Study